OpenStudy (anonymous):

how do you get the value of csc (5pi/4) with the unit circle

OpenStudy (mathstudent55):

First, it helps to know this: \( \csc x= \dfrac{1}{\sin x} \)

OpenStudy (jdoe0001):

you use a "good" unit circle, a good one will have it there :)

OpenStudy (jdoe0001):

oh, ohh csc, well csc = 1/sin as mathstudent55 said, so your "good" unit circle will have the sine functions too :)

OpenStudy (mathstudent55):